# On the non-archimedean Monge–Ampère equation in mixed characteristic

### v2

@@ 5. Continuity of the envelope of the zero function
Let $K$ be a complete discretely valued field of mixed characteristic $(0,p)$. Let $X$ be a $n$-dimensional smooth projective variety over $K$ and $L$ an ample line bundle on $X$. In the approach of Boucksom–Favre–Jonsson [BFJ16, BFJ15] to pluripotential theory on $X^{\rm an}$, a model ${\mathscr{L}}$ of $L$ on the model ${\mathscr{X}}$ induces a curvature form $\theta$ on $X^{\rm an}$ and we denote by $P_{\theta}(f)$ the $\theta$-psh envelope of a continuous real function $f$ on $X^{\rm an}$, see [GJKM19, 2.4–2.6] for details. We follow the strategy from [BFJ16] and [GJKM19] to show continuity of the envelope of the zero function $P_{\theta}(0)$.
@@
is globally generated. Taking the tensor product of (5.1) and (5.2) we see that $\mathfrak{b}_{m}\otimes{\mathscr{A}}^{\otimes m_{0}}\otimes{\mathscr{L}}^{\otimes m}$ is globally generated. Finally (a), (b) and (c) imply our claim by the strategy of the proof of [BFJ16, Thm. 8.5]. For convenience of the reader, we give here some details. For $m\gg 0$, let
@@
which is a super-additive sequence of $\theta$-psh functions as shown at the beginning of the proof of loc. cit.. Then Step 1 of the quoted proof shows that $P_{\theta}(0)=\sup_{m}\varphi_{m}=\lim_{m}\varphi_{m}$ on the quasi-monomial points of ${X^{\mathrm{an}}}$. Using [GJKM19, Proposition 2.10], this holds pointwise on the whole ${X^{\mathrm{an}}}$. We also consider the functions
@@ 6. Resolution of singularities
The arguments are the same as for [GJKM19, Lemma 7.5]. In a first step, we start with a projective model ${\mathscr{Y}}$ of $X$ such that $L$ extends to an ample line bundle ${\mathscr{H}}$ on ${\mathscr{Y}}$, possibly replacing $L$ by a positive tensor power. By a result of Lütkebohmert [Lüt93, Lemma 2.2], there is a blowing up $\pi\colon\mathscr{Z}\to{\mathscr{Y}}$ in an ideal sheaf supported in the special fiber of ${\mathscr{Y}}$ such that the identity on $X$ extends to a morphism $\mathscr{Z}\to{\mathscr{X}}$. A property of blowing ups [Har77, Proposition II.7.13] shows that $\pi^{*}({\mathscr{H}}^{\otimes\ell})\otimes{\mathscr{O}}_{\mathscr{Z}/{\mathscr{Y}}}(1)$ is an ample line bundle on $\mathscr{Z}$ with generic fiber $L^{\otimes\ell}$ for sufficiently large $\ell$. Replacing ${\mathscr{X}}$ by $\mathscr{Z}$ and $L$ by $L^{\otimes\ell}$, we conclude that we may assume that $L$ extends to an ample line bundle ${\mathscr{H}}$ on ${\mathscr{X}}$. By a result of Pépin [Pép13, Thm. 3.1], there is a a blowing-up morphism $\pi^{\prime}\colon\mathscr{Z}^{\prime}\to{\mathscr{X}}$ centered in the special fiber of ${\mathscr{X}}$ such that $\mathscr{Z}^{\prime}$ is semi-factorial. So similarly as in the first step, replacing ${\mathscr{X}}$ by $\mathscr{Z}^{\prime}$ and $L$ by a positive tensor power, we may assume that $L$ extends to an ample line bundle ${\mathscr{H}}$ on a semi-factorial projective model ${\mathscr{X}}$. This is the conclusion of the second step. Then we apply resolution of singularities to ${\mathscr{X}}$ to get a projective $S$-morphism $\pi\colon{\mathscr{X}}^{\prime}\to{\mathscr{X}}$ which is an isomorphism on generic fibers and so we may identify the generic fiber of ${\mathscr{X}}^{\prime}$ with $X$. Since $\pi$ is projective, there is an $\ell>0$ such that ${\mathscr{L}}^{\prime}\coloneqq\pi^{*}({\mathscr{H}}^{\otimes\ell})\otimes{\mathscr{O}}_{{\mathscr{X}}^{\prime}/{\mathscr{X}}}(1)$ is an ample line bundle on ${\mathscr{X}}^{\prime}$. However, we do not know if $\pi$ is a blow up in an ideal supported in the special fiber and hence the restriction $F$ of ${\mathscr{O}}_{{\mathscr{X}}^{\prime}/{\mathscr{X}}}(1)$ to the generic fiber $X$ of ${\mathscr{X}}^{\prime}$ might be non-trivial. Since ${\mathscr{X}}$ is semi-factorial, $F$ extends to a line bundle ${\mathscr{F}}$ on ${\mathscr{X}}$ and we can replace ${\mathscr{O}}_{{\mathscr{X}}^{\prime}/{\mathscr{X}}}(1)$ by ${\mathscr{O}}_{{\mathscr{X}}^{\prime}/{\mathscr{X}}}(1)\otimes\varphi^{*}({\mathscr{F}}^{-1})$. Then ${\mathscr{L}}^{\prime}$ is a model of $L^{\otimes\ell}$ proving the claim. ∎
@@ 8. The Monge–Ampère equation
Uniqueness follows from a result of Yuan and Zhang, see [BFJ15, §8.1]. To prove existence of a $\theta$-psh solution $\varphi$, the variational method of Boucksom, Favre, and Jonsson is used. Continuity of the envelope for ample line bundles on $X$ holds by Theorem 7.1. By [${}^{+}$20, Theorems 6.3.2, 6.3.3], we conclude that $\theta$ satisfies the orthogonality property. Then existence of a continuous solutions follows from [BFJ15, Theorem 8.1]. ∎
@@ Acknowledgements
We are grateful to Bhargav Bhatt, Huayi Chen and Mattias Jonsson for remarks about previous versions of this paper. We are especially grateful to Karl Schwede for discussions and help about subadditivity of global $+$-test ideals and perturbation friendly global test ideals.

### v3

@@ 5. Continuity of the envelope of the zero function
Let $K$ be a complete discretely valued field of mixed characteristic $(0,p)$. Let $X$ be a $n$-dimensional smooth projective variety over $K$ and $L$ an ample line bundle on $X$. In the approach of Boucksom–Favre–Jonsson [BFJ16, BFJ15] to pluripotential theory on $X^{\rm an}$, a model $\mathscr{L}$} of $L$ on the model $\mathscr{X}$} induces a curvature form $\theta$ on $X^{\rm an}$ and we denote by $P_{\theta}(u)$ the $\theta$-psh envelope of a continuous real function $u$ on $X^{\rm an}$, see [GJKM19, 2.4–2.6] for details. We follow the strategy from [BFJ16] and [GJKM19] to show continuity of the envelope of the zero function $P_{\theta}(0)$.
@@
is globally generated. Taking the tensor product of (5.1) and (5.2) we see that $\mathfrak{b}_{m}\otimes{\mathscr{A}}^{\otimes m_{0}}\otimes{\mathscr{L}}^{\otimes m}$ is globally generated.
Finally (a), (b) and (c) imply our claim by the strategy of the proof of [BFJ16, Thm. 8.5]. For convenience of the reader, we give here some details. For $m\gg 0$, let
@@
which is a super-additive sequence of $\theta$-psh functions as shown at the beginning of the proof of loc. cit.. Then Step 1 of the quoted proof shows that $P_{\theta}(0)=\sup_{m}\varphi_{m}=\lim_{m}\varphi_{m}$ on the quasi-monomial points of $X^{\mathrm{an}}$}. Using [GJKM19, Proposition 2.10], this holds pointwise on the whole $X^{\mathrm{an}}$}.
We also consider the functions
@@ 6. Resolution of singularities
The arguments are the same as for [GJKM19, Lemma 7.5]. In a first step, we start with a projective model $\mathscr{Y}$} of $X$ such that $L$ extends to an ample line bundle $\mathscr{H}$} on $\mathscr{Y}$}, possibly replacing $L$ by a positive tensor power. By a result of Lütkebohmert [Lüt93, Lemma 2.2], there is a blowing up $\pi\colon\mathscr{Z}\to{\mathscr{Y}}$ in an ideal sheaf supported in the special fiber of $\mathscr{Y}$} such that the identity on $X$ extends to a morphism $\mathscr{Z}\to{\mathscr{X}}$. A property of blowing ups [Har77, Proposition II.7.13] shows that $\pi^{*}({\mathscr{H}}^{\otimes\ell})\otimes{\mathscr{O}}_{\mathscr{Z}/{\mathscr{Y}}}(1)$ is an ample line bundle on $\mathscr{Z}$ with generic fiber $L^{\otimes\ell}$ for sufficiently large $\ell$. Replacing $\mathscr{X}$} by $\mathscr{Z}$ and $L$ by $L^{\otimes\ell}$, we conclude that we may assume that $L$ extends to an ample line bundle $\mathscr{H}$} on $\mathscr{X}$}.
By a result of Pépin [Pép13, Thm. 3.1], there is a a blowing-up morphism $\pi^{\prime}\colon\mathscr{Z}^{\prime}\to{\mathscr{X}}$ centered in the special fiber of $\mathscr{X}$} such that $\mathscr{Z}^{\prime}$ is semi-factorial. So similarly as in the first step, replacing $\mathscr{X}$} by $\mathscr{Z}^{\prime}$ and $L$ by a positive tensor power, we may assume that $L$ extends to an ample line bundle $\mathscr{H}$} on a semi-factorial projective model $\mathscr{X}$}. This is the conclusion of the second step.
Then we apply resolution of singularities to $\mathscr{X}$} to get a projective $S$-morphism $\pi\colon{\mathscr{X}}^{\prime}\to{\mathscr{X}}$ which is an isomorphism on generic fibers and so we may identify the generic fiber of $\mathscr{X}$$^{\prime}$ with $X$. Since $\pi$ is projective, there is an $\ell>0$ such that ${\mathscr{L}}^{\prime}\coloneqq\pi^{*}({\mathscr{H}}^{\otimes\ell})\otimes{\mathscr{O}}_{{\mathscr{X}}^{\prime}/{\mathscr{X}}}(1)$ is an ample line bundle on $\mathscr{X}$$^{\prime}$. However, we do not know if $\pi$ is a blow up in an ideal supported in the special fiber and hence the restriction $F$ of ${\mathscr{O}}_{{\mathscr{X}}^{\prime}/{\mathscr{X}}}(1)$ to the generic fiber $X$ of $\mathscr{X}$$^{\prime}$ might be non-trivial. Since $\mathscr{X}$} is semi-factorial, $F$ extends to a line bundle $\mathscr{F}$} on $\mathscr{X}$} and we can replace ${\mathscr{O}}_{{\mathscr{X}}^{\prime}/{\mathscr{X}}}(1)$ by ${\mathscr{O}}_{{\mathscr{X}}^{\prime}/{\mathscr{X}}}(1)\otimes\varphi^{*}({\mathscr{F}}^{-1})$. Then $\mathscr{L}$$^{\prime}$ is a model of $L^{\otimes\ell}$ proving the claim. ∎
@@ 8. The Monge–Ampère equation
Uniqueness follows from a result of Yuan and Zhang, see [BFJ15, §8.1]. To prove existence of a $\theta$-psh solution $\varphi$, the variational method of Boucksom, Favre, and Jonsson is used. Continuity of the envelope for ample line bundles on $X$ holds by Theorem 7.1. By [BGJ+20, Theorems 6.3.2, 6.3.3], we conclude that $\theta$ satisfies the crucial orthogonality property (see [BFJ15, Definition 7.1] for the definition). Then existence of a continuous solution follows from [BFJ15, Theorem 8.1]. ∎
@@ Acknowledgements
We are grateful to Bhargav Bhatt, Huayi Chen and Mattias Jonsson for remarks about previous versions of this paper. We are especially grateful to Karl Schwede for discussions and help about subadditivity of global ++-test ideals and perturbation friendly global test ideals. We are grateful to the referee for helpful remarks and suggestions.